OpenStudy (anonymous):

if cos(n*pi)=(-1)^n, sin(n*pi)=0 what is cos(n*pi/2) sin(n*pi/2) sin(n+1)*pi/2 cos(n+1)*pi/2 sin(n-1)*pi/2 cos(n-1)*pi/2 sin(n-1)*pi cos(n-1)*pi

OpenStudy (anonymous):

sin(n*pi)=0 sin(n+1)*pi/2=(-1)^n for n=odd >>0 cos(n+1)*pi/2=0 sin(n-1)*pi/2=(-1)^(n-1) for n=odd >>0 cos(n-1)*pi/2 =0 sin(n-1)*pi=0 cos(n-1)*pi=(-1)^(n-1)

OpenStudy (dumbcow):

cos(n*pi/2) = (i^n)[1+(-1)^n]/2

OpenStudy (dumbcow):

sin(n*pi/2) = (i^n-1)[1-(-1)^n]/2